# Cointegration with Occasionally Binding Constraints

### v3

@@ 4.4 Deterministic trends
Since we would only apply the CKSVAR when $\{y_{t}\}$ spent an appreciable portion of the sample on both sides of zero, it is appropriate to work within an asymptotic framework in which both regions continue to be visited with nonvanishing probability as $T\rightarrow\infty$. [^14] Here we maintain that the location of the threshold delimiting the regimes is normalised to zero; of course the argument continues to apply if that threshold were instead fixed at some other finite level. The cleanest way to ensure this is to restrict $c$ such that a deterministic trend may be imparted to $x_{t}$, but not to $y_{t}$. In the context of cases (i) and (ii), essentially the same condition is required, but is denoted slightly differently so as to be consistent with the notation used in the analysis of these two cases. (Recall the definitions of $P_{\beta_{\perp}^{+}}$ and $P_{\beta_{\perp}}(+1)$ appearing in (LABEL:Pbetaplus) and (LABEL:projk1) above.)
@@
Suppose that DGP, ERR, DET and CVAR hold. If additionally CO(i) (respectively CO(ii)) holds, then the conclusions of 4.1 (respectively 4.2) hold with with $X_{n}^{d}$ in place of $X_{n}$.
@@
(i). The preceding illustrates how common stochastic and deterministic trends may be simultaneously accomodated within a CKSVAR. In particular, because the deterministic trends enter $z_{t}$ through $(P_{\beta_{\perp}^{+}}c)t$ (or $[P_{\beta_{\perp}}(+1)c]t$), it remains true that $\beta^{+\mathsf{T}}z_{t}\sim I^{\ast}(0)$ in case (i) and $\beta(y_{t})^{\mathsf{T}}z_{t}\sim I^{\ast}(0)$ in case (ii), since these transformations of $z_{t}$ now eliminate both its common stochastic and deterministic trends. These objects thus continue to describe the (nonlinear) cointegrating relations between the elements of $z_{t}$.
@@ 5 Conclusion
Our results provide a complete characterisation of the forms of nonlinear cointegration (between processes $\ast$-integrated of order one) generated by the CKSVAR. In deriving these, we have given the first treatment of how nonlinear cointegration, in the profound sense of nonlinear common stochastic trends and nonlinear cointegrating relations, may be systematically generated within a (nonlinear) VAR, and thus the first extension of the Granger–Johansen representation theorem to a nonlinearly cointegrated setting. The special structure of the CKSVAR makes this problem peculiarly tractable, while being flexible enough to generate interesting departures from linear cointegration. Our results indicate how progress may now be made in the analysis of more general nonlinear VARs with unit roots, while our representation theory provides the foundations for inference on cointegrating relations in the CKSVAR. Our findings with respect to these problems will be reported elsewhere; some initial results regarding inference in this setting, in the univariate case, are given in Bykhovskaya and Duffy (2024).
@@ Appendix A Auxiliary lemmas
We here collect three elementary lemmas, whose proofs appear in C for completeness. The third records, for convenience, some well-known properties of the linear cointegrated VAR.
@@
Then there exists a $C<\infty$ such that for all $n\in\mathbb{N}$ and $m\in[1,m_{0}]$,
@@
Suppose $\Phi(\lambda)\coloneqq I_{p}-\sum_{i=1}^{k}\Phi_{i}\lambda^{i}$ and $\Pi\coloneqq-\Phi(1)$ satisfy CVAR (i.e. without the ‘$\pm$’ superscripts). Define
@@
and let $\Gamma(\lambda)\coloneqq I_{p}-\sum_{i=1}^{k-1}\Gamma_{i}\lambda^{i}$ be such that $\Phi(\lambda)=\Phi(1)\lambda+\Gamma(\lambda)(1-\lambda)$. [^15] As per the footnote to (LABEL:betaplus), the definition of $\boldsymbol{\Pi}$ here should be read ‘row-wise’, so that for $m\in\{2,\ldots,k\}$, rows $(m-1)p+1$ to $mp$ of $\boldsymbol{\Pi}$ are given by $[0_{p\times(m-2)p},I_{p},-I_{p},0_{p\times(k-m)p}]$. The matrix definitions appearing subsequently in these appendices should be interpreted similarly. Then:
@@ Appendix B Proofs of theorems
and recalling $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$, we may rewrite the model (LABEL:vecm) as
@@
so that $\boldsymbol{\Pi}^{+}=\boldsymbol{\alpha}^{+}\boldsymbol{\beta}^{+\mathsf{T}}$ (by A.3), where $\boldsymbol{\alpha}^{+}$ and $\boldsymbol{\beta}^{+}$ are as in (LABEL:betaplus), we render the system in companion form as
@@
[…]
@@
where $\boldsymbol{\beta}_{1:p}^{+}$ denotes the first $p$ rows of $\boldsymbol{\beta}^{+}$. Recall the definition of $\boldsymbol{F}_{\delta}$ given in (LABEL:Fkgeneral) above.
@@
$\delta_{t}$ =$=\begin{cases}(y_{t-1}^{+}+\text{$\overline{y}$}_{t})/\text{$\overline{y}$}_{t}&\text{if }y_{t-1}^{+}+\text{$\overline{y}$}_{t}<0,\\ 0&\text{otherwise};\end{cases}$ (B.4)
@@
Premultiply (LABEL:poscompanion) by $\boldsymbol{\beta}^{+\mathsf{T}}$ and then use (LABEL:vdef) to obtain
@@
To obtain the law of motion for $y_{t}^{-}$, we take the first equation in (LABEL:poscompanion)
@@
(where $e_{1}$ and $\boldsymbol{e}_{1}$ denote the first column of $I_{p}$ and $I_{kp}$ respectively), which using (LABEL:vdef) can be rearranged as
@@
[…]
@@
where $\delta_{t}\in[0,1]$ is defined in (LABEL:delcens). (Observe that if $y_{t-1}^{+}+\text{$\overline{y}$}_{t}<0$, then $\overline{y}_{t}<-y_{t-1}^{+}\leq 0$, so this object is well defined.) Making the substitution $y_{t-1}^{-}=\delta_{t-1}\text{$\overline{y}$}_{t-1}$ on the r.h.s. of (LABEL:xirep) and (LABEL:pseudy), we see that the trajectories of $\{\boldsymbol{\xi}_{t}^{+}\}$ and $\{y_{t}^{-}\}$ for $t\geq 1$ (conditional on the initial values $\{z_{t}\}_{t=-k+1}^{0}$) will be reproduced exactly by
@@
[…]
@@
with $\delta_{0}\coloneqq 1$ and $\overline{y}_{0}\coloneqq y_{0}^{-}$. Finally, we note that (LABEL:srcomp) provides the companion form representation of this system. ∎
By Jungers (2009, Prop. 1.8), CO(i).2 implies that $\rho_{{\scriptstyle\mathrm{JSR}}}(\{\boldsymbol{F}_{\delta}\mid\delta\in[0,1]\})=\rho_{{\scriptstyle\mathrm{JSR}}}(\{\boldsymbol{F}_{0},\boldsymbol{F}_{1}\})<1$. In view of ERR and CO(i).4, applying A.1 to the representation (LABEL:srcomp) then yields that $\sup_{t\in\mathbb{N}}\lVert\boldsymbol{\zeta}_{t}\rVert_{2+\delta_{u}}<\infty$, and hence $\boldsymbol{\zeta}_{t}\sim I^{\ast}(0)$. In particular, $\boldsymbol{\xi}_{t}^{+}\sim I^{\ast}(0)$ and $y_{t}^{-}\sim I^{\ast}(0)$, which gives part (i) of the theorem.
@@
for $P_{\beta_{\perp}^{+}}$ as defined in (LABEL:Pbetaplus), noting only at the very end of the proof how the result simplifies when DET holds. As discussed in 4.4, the condition (LABEL:detprime-case1) (which appears there as DET) permits the model to impart a deterministic trend to $x_{t}$, but not to $y_{t}$.
@@
(see A.3). Premutiplying (LABEL:poscompanion) by $\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}$ and cumulating yields
@@
Since the top left $p\times p$ block of $\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}$ is $P_{\beta_{\perp}^{+}}=\beta_{\perp}^{+}[\alpha_{\perp}^{+\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}]^{-1}\alpha_{\perp}^{+\mathsf{T}}$ (A.3), the preceding implies
@@
[…]
where the final equality follows since ${\cal Z}_{0}\in\mathscr{M}\subset\operatorname{sp}\beta_{\perp}^{+}$. Now recalling (LABEL:vdef),
@@
we have from (LABEL:ztsmldecmp) that
@@
Under (LABEL:detprime-case1), the first equation in (LABEL:reginit) is
@@
[…]
@@
on $D[0,1]$, where the penultimate equality follows from $P_{\beta_{\perp}^{+}}\Gamma^{+}(1){\cal Z}_{0}={\cal Z}_{0}$, as per (LABEL:cocens-init) above. Therefore by Bykhovskaya and Duffy (2024, Lem. A.1), and $y_{t}^{-}\sim I^{\ast}(0)$,
@@
Since (LABEL:detprime-case1) permits a deterministic trend to be present in $x_{t}$, as evident from (LABEL:reginit) above, we consider the linearly detrended process
@@
where $E_{-1}\in\mathbb{R}^{p\times(p-1)}$ collects the final $p-1$ columns of $I_{p}$. To obtain the weak limit of $n^{-1/2}x_{\lfloor n\lambda\rfloor}^{d},$ we substitute
@@
from (LABEL:reginit1) into the final $p-1$ equations from (LABEL:reginit), to obtain
@@
for $E_{-1}$ the final $p-1$ columns of $I_{p}$. Hence, by $\eta_{t}\sim I^{\ast}(0)$, ERR, (LABEL:cocens-init) and (LABEL:ycenslim),
@@
Finally, we suppose that DET holds in place of (LABEL:detprime-case1). In this case, $P_{\beta_{\perp}^{+}}c=0$ and therefore $x_{t}^{d}=x_{t}$ for all $t\in\{1,\ldots,T\}$. Thus part (ii) of the theorem follows from (LABEL:ycenslim) and (LABEL:xcenslim). ∎
B.2 Proof of 4.2
Because $\{y_{t}\}$ now behaves like an integrated process in both the positive and negative regimes, we have to take quite a different approach from that utilised in case (i). Recall $z_{t}^{\mathsf{T}}=(y_{t},x_{t}^{\mathsf{T}})$, but now also define $z_{t}^{\ast\mathsf{T}}\coloneqq(y_{t}^{+},y_{t}^{-},x_{t}^{\mathsf{T}})$, so that (LABEL:vecm) may be written as
@@
[…]
so that $\boldsymbol{\Pi}(y)=\boldsymbol{\alpha}\boldsymbol{\beta}(y)^{\mathsf{T}}$, where $\boldsymbol{\alpha}$, $\boldsymbol{\beta}(y)$ and $S(y)$ are as defined in (LABEL:alpabeta)–(LABEL:S), we can rewrite the system in companion form as
@@
Our next result is the counterpart of B.1 for the present case, in that it provides a nonlinear VAR representation for the short-memory components of the model: the equilibrium errors $\xi_{t}\coloneqq\beta(y_{t})^{\mathsf{T}}z_{t}$ and the differences $\Delta z_{t}^{\ast}$, as collected in
@@
Suppose that DGP, CVAR and CO(i).1 hold. Let
@@
Using the definition of $\boldsymbol{\xi}_{t}$, we may write (LABEL:vecmcoclas) as
@@
Let $\Delta\boldsymbol{\beta}(y_{t})\coloneqq\boldsymbol{\beta}(y_{t})-\boldsymbol{\beta}(y_{t-1})$, and note that $[\Delta\boldsymbol{\beta}(y_{t})]^{\mathsf{T}}\boldsymbol{z}_{t}=[\Delta\boldsymbol{\beta}(y_{t})]_{1:1}^{\mathsf{T}}y_{t}$, for $[\Delta\boldsymbol{\beta}(y_{t})]_{1:1}$ the first row of $\Delta\boldsymbol{\beta}(y_{t})$, since all its other rows are zero. Thus with the aid of (LABEL:delzxi) we obtain
@@
By construction, for $\delta_{t}$ as defined in (LABEL:delocens), we have that $y_{t}=\delta_{t}\Delta y_{t}$ whenever $y_{t-1}$ and $y_{t}$ have opposite signs. Since this is also the only case in which $\Delta\boldsymbol{\beta}(y_{t})\neq 0$, it follows that
@@
Substituting this into (LABEL:xirepcase2) and recalling the definition of $\overline{\boldsymbol{\beta}}_{t}$ thus yields (LABEL:xicase2).
@@
Since $\overline{\boldsymbol{\beta}}_{t}$ is always a convex combination of $\boldsymbol{\beta}(-1)$ and $\boldsymbol{\beta}(+1)$, the autoregressive matrices in the representation (LABEL:xicase2) are contained in the set
@@
By Jungers (2009, Prop. 1.8), $\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A})=\rho_{{\scriptstyle\mathrm{JSR}}}(\{I+\boldsymbol{\beta}(+1)^{\mathsf{T}}\boldsymbol{\alpha},I+\boldsymbol{\beta}(-1)^{\mathsf{T}}\boldsymbol{\alpha}\})$, which is strictly less than unity by CO(ii).2. That $\boldsymbol{\xi}_{t}\sim I^{\ast}(0)$ now follows from ERR, CO(ii).4 and A.1. In particular, $\xi_{t}=\beta(y_{t})^{\mathsf{T}}z_{t}$, $\Delta y_{t}^{+}$, $\Delta y_{t}^{-}$ and $\Delta x_{t}$ are all $I^{\ast}(0)$, and hence so too is
@@
For the purposes of proving part (ii) of the theorem, we shall initially suppose (as we did in the proof of 4.1) that DET is replaced by the weaker condition
@@
for $P_{\beta_{\perp}}(+1)$ as defined in (LABEL:projk1), noting only at the very end of the proof how the result simplifies when DET holds.
Our next step is to extract the common trend components. Recall the definitions of $\beta_{\perp}(y)$ and $\Gamma(1;y)$ in (LABEL:betaperpfn), and note that we may take
@@
where the r.h.s. is invertible for each $y\in\mathbb{R}$ by A.3; it follows that the complementary projections
@@
are also well defined for each $y\in\mathbb{R}$.
Now premultiplying (LABEL:vecmcoclas) by $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}$ and cumulating yields
@@
[…]
Hence, using the result of part (i), the first $p$ rows of (LABEL:z-costat-proj) give

### v4

@@ 4.4 Deterministic trends
Since we would only apply the CKSVAR when $\{y_{t}\}$ spent an appreciable portion of the sample on both sides of zero, it is appropriate to work within an asymptotic framework in which both regions continue to be visited with nonvanishing probability as $T\rightarrow\infty$. [^13] Here we maintain that the location of the threshold delimiting the regimes is normalised to zero; of course the argument continues to apply if that threshold were instead fixed at some other finite level. The cleanest way to ensure this is to restrict $c$ such that a deterministic trend may be imparted to $x_{t}$, but not to $y_{t}$. In the context of cases (i) and (ii), essentially the same condition is required, but is denoted slightly differently so as to be consistent with the notation used in the analysis of these two cases. (Recall the definitions of $P_{\beta_{\perp}^{+}}$ and $P_{\beta_{\perp}}(+1)$ appearing in (4.3) and (4.7) above.)
@@
Suppose that DGP, ERR, DET and CVAR hold. If additionally CO(i) (respectively CO(ii)) holds, then the conclusions of Theorem 4.1 (respectively Theorem 4.2) hold with with $X_{n}^{d}$ in place of $X_{n}$.
@@
(i). The preceding illustrates how common stochastic and deterministic trends may be simultaneously accomodated within a CKSVAR. In particular, because the deterministic trends enter $z_{t}$ through $(P_{\beta_{\perp}^{+}}c)t$ (or $P_{\beta_{\perp}}(+1)c]t$), it remains true that $\beta^{+\mathsf{T}}z_{t}\sim I^{\ast}(0)$ in case (i) and $\beta(y_{t})^{\mathsf{T}}z_{t}\sim I^{\ast}(0)$ in case (ii), since these transformations of $z_{t}$ now eliminate both its common stochastic and deterministic trends. These objects thus continue to describe the (nonlinear) cointegrating relations between the elements of $z_{t}$.
@@ 5 Conclusion
Our results provide a complete characterisation of the forms of nonlinear cointegration (between processes $\ast$-integrated of order one) generated by the CKSVAR. In deriving these, we have given the first treatment of how nonlinear cointegration, in the profound sense of nonlinear common stochastic trends and nonlinear cointegrating relations, may be systematically generated within a (nonlinear) VAR, and thus the first extension of the Granger–Johansen representation theorem to a nonlinearly cointegrated setting. The special structure of the CKSVAR makes this problem peculiarly tractable, while being flexible enough to generate interesting departures from linear cointegration. Our results indicate how progress may now be made in the analysis of more general nonlinear VARs with unit roots, while our representation theory provides the foundations for inference on cointegrating relations in the CKSVAR. Our findings with respect to these problems will be reported elsewhere; some initial results regarding inference in this setting, in the univariate case, are given in Bykhovskaya and Duffy (2024) and Duffy and Jiao (2025).
@@ Appendix A Auxiliary lemmas
We here collect three elementary lemmas, whose proofs appear in Appendix C for completeness. The third records, for convenience, some properties of the linear cointegrated VAR which are closely related to Lemmas A.1 and A.2 in Hansen (2005).
@@
Then for each $\gamma\in(\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A}),1)$, there exists a $C<\infty$ such that for all $t\in\mathbb{N}$,
$\lVert w_{t}\rVert\leq C\left[\sum_{s=0}^{t-1}\gamma^{s}(1+\lVert v_{t-s}\rVert)+\gamma^{t}\lVert w_{0}\rVert\right],$
and for all $n\in\mathbb{N}$ and $m\in[1,m_{0}]$,
@@
Suppose $\Phi(\lambda)\coloneqq I_{p}-\sum_{i=1}^{k}\Phi_{i}\lambda^{i}$ and $\Pi\coloneqq-\Phi(1)$ satisfy CVAR (i.e. without the $\pm$’ superscripts). Define
@@
and let $\Gamma(\lambda)\coloneqq I_{p}-\sum_{i=1}^{k-1}\Gamma_{i}\lambda^{i}$ be such that $\Phi(\lambda)=\Phi(1)\lambda+\Gamma(\lambda)(1-\lambda)$. [^14] As per the footnote to (4.2), the definition of $\boldsymbol{\Pi}$ here should be read ‘row-wise’, so that for $m\in\{2,\ldots,k\}$, rows (m1)p+1(m-1)p+1 to $mp$ of $\boldsymbol{\Pi}$ are given by $0_{p\times(m-2)p},I_{p},-I_{p},0_{p\times(k-m)p}]$. The matrix definitions appearing subsequently in these appendices should be interpreted similarly. Then:
@@ Appendix B Proofs of theorems
and recalling $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$, we may rewrite the model (3.2) as
@@
so that $\boldsymbol{\Pi}^{+}=\boldsymbol{\alpha}^{+}\boldsymbol{\beta}^{+\mathsf{T}}$, where $\boldsymbol{\alpha}^{+}$ and $\boldsymbol{\beta}^{+}$ are as in (4.2), we render the system in companion form as
@@
where $\xi_{t}^{+}\coloneqq\beta^{+\mathsf{T}}z_{t}$; (b) the lagged levels $\boldsymbol{y}_{t-1}^{-}\coloneqq(y_{t-1}^{-},\ldots,y_{t-k+1}^{-})^{\mathsf{T}}$; and (c) an auxiliary series $\text{$\overline{y}$}_{t}$. Collect these in $\boldsymbol{\zeta}_{t}\in\mathbb{R}^{p(k-1)+k+r}$, and conformably define
@@
where $\boldsymbol{\beta}_{1:p}^{+}$ denotes the first $p$ rows of $\boldsymbol{\beta}^{+}$. Recall the definition of $\boldsymbol{F}_{\delta}$ given in (4.1) above.
@@
$\delta_{t}$ $=\begin{cases}(y_{t-1}^{+}+\text{$\overline{y}$}_{t})/\text{$\overline{y}$}_{t}&\text{if }y_{t-1}^{+}+\text{$\overline{y}$}_{t}<0,\\ 0&\text{otherwise};\end{cases}$ (B.4)
@@
Premultiply (B.2) by $\boldsymbol{\beta}^{+\mathsf{T}}$ and then use (B.1) to obtain
@@
To obtain the law of motion for $y_{t}^{-}$, we take the first equation in (B.2)
@@
(where $e_{1}$ and $\boldsymbol{e}_{1}$ denote the first column of $I_{p}$ and $I_{kp}$ respectively), which using (B.1) can be rearranged as
@@
It follows that if we define $\text{$\overline{y}$}_{t}$ to equal the r.h.s., then $y_{t}=y_{t-1}^{+}+\text{$\overline{y}$}_{t}$, and so
@@
where $\delta_{t}\in[0,1]$ is defined in (B.4). (Observe that if $y_{t-1}^{+}+\text{$\overline{y}$}_{t}<0$, then $\overline{y}_{t}<-y_{t-1}^{+}\leq 0$, so this object is well defined.) Making the substitution $y_{t-1}^{-}=\delta_{t-1}\text{$\overline{y}$}_{t-1}$ on the r.h.s. of (B.5) and (B.6), we see that the trajectories of $\{\boldsymbol{\xi}_{t}^{+}\}$ and $\{y_{t}^{-}\}$ for $t\geq 1$ (conditional on the initial values $\{z_{t}\}_{t=-k+1}^{0}$) will be reproduced exactly by
@@
$\text{$\overline{y}$}_{t}$ $=e_{1}^{\mathsf{T}}c+[1+e_{1}^{\mathsf{T}}(\phi_{1}^{-}-\phi_{1}^{+})]\delta_{t-1}\text{$\overline{y}$}_{t-1}$
@@
with $\delta_{0}\coloneqq 1$ and $\overline{y}_{0}\coloneqq y_{0}^{-}$. Finally, we note that (B.3) provides the companion form representation of this system. ∎
By Jungers (2009, Prop. 1.8), CO(i).2 implies that $\rho_{{\scriptstyle\mathrm{JSR}}}(\{\boldsymbol{F}_{\delta}\mid\delta\in[0,1]\})=\rho_{{\scriptstyle\mathrm{JSR}}}(\{\boldsymbol{F}_{0},\boldsymbol{F}_{1}\})<1$. In view of ERR and CO(i).4, applying Lemma A.1 to the representation (B.3) then yields that $\sup_{t\in\mathbb{N}}\lVert\boldsymbol{\zeta}_{t}\rVert_{2+\delta_{u}}<\infty$, and hence $\boldsymbol{\zeta}_{t}\sim I^{\ast}(0)$. In particular, $\boldsymbol{\xi}_{t}^{+}\sim I^{\ast}(0)$ and $y_{t}^{-}\sim I^{\ast}(0)$, which gives part (i) of the theorem.
@@
for $P_{\beta_{\perp}^{+}}$ as defined in (4.3), noting only at the very end of the proof how the result simplifies when DET holds. As discussed in Section 4.4, the condition (B.8) (which appears there as DET) permits the model to impart a deterministic trend to $x_{t}$, but not to $y_{t}$.
@@
(see Lemma A.3). Premutiplying (B.2) by $\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}$ and cumulating yields
@@
Since the top left $p\times p$ block of $\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}$ is $P_{\beta_{\perp}^{+}}=\beta_{\perp}^{+}[\alpha_{\perp}^{+\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}]^{-1}\alpha_{\perp}^{+\mathsf{T}}$ (Lemma A.3), the preceding implies
@@
where $z_{\xi,t}\coloneqq[\boldsymbol{z}_{\xi,t}]_{1:p}$. Under CO(i).4, $n^{-1/2}z_{t}\overset{p}{\rightarrow}{\cal Z}_{0}$ for each $t\in\{-k+1,\ldots,0\}$; hence
$n^{-1/2}[\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\boldsymbol{z}_{0}]_{1:p}\overset{p}{\rightarrow}\beta_{\perp}^{+}[\alpha_{\perp}^{+\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}]^{-1}\alpha_{\perp}^{+\mathsf{T}}\Gamma^{+}(1){\cal Z}_{0}={\cal Z}_{0}$ (B.10)
where the final equality follows since ${\cal Z}_{0}\in\mathscr{M}\subset\operatorname{sp}\beta_{\perp}^{+}$. Now recalling (B.1),
@@
we have from (B.9) that
@@
Under (B.8), the first equation in (B.11) is
@@
As noted above $\eta_{t}\sim I^{\ast}(0)$, while $n^{-1/2}y_{0}\overset{p}{\rightarrow}{\cal Y}_{0}$ by CO(i).4. Hence by ERR,
@@
on $D[0,1]$, where the penultimate equality follows from $P_{\beta_{\perp}^{+}}\Gamma^{+}(1){\cal Z}_{0}={\cal Z}_{0}$, as per (B.10) above. Therefore by Bykhovskaya and Duffy (2024, Lem. A.1), and $y_{t}^{-}\sim I^{\ast}(0)$,
@@
Since (B.8) permits a deterministic trend to be present in $x_{t}$, as evident from (B.11) above, we consider the linearly detrended process
@@
where $E_{-1}\in\mathbb{R}^{p\times(p-1)}$ collects the final p1p-1 columns of $I_{p}$. To obtain the weak limit of $n^{-1/2}x_{\lfloor n\lambda\rfloor}^{d},$ we substitute
@@
from (B.12) into the final p1p-1 equations from (B.11), to obtain
@@
for $E_{-1}$ the final p1p-1 columns of $I_{p}$. Hence, by $\eta_{t}\sim I^{\ast}(0)$, ERR, (B.10) and (B.13),
@@
Finally, we suppose that DET holds in place of (B.8). In this case, $P_{\beta_{\perp}^{+}}c=0$ and therefore $x_{t}^{d}=x_{t}$ for all $t\in\{1,\ldots,T\}$. Thus part (ii) of the theorem follows from (B.13) and (B.14). ∎
B.2 Proof of Theorem 4.2
Because $\{y_{t}\}$ now behaves like an integrated process in both the positive and negative regimes, we have to take quite a different approach from that utilised in case (i). Recall $z_{t}^{\mathsf{T}}=(y_{t},x_{t}^{\mathsf{T}})$, but now also define $z_{t}^{\ast\mathsf{T}}\coloneqq(y_{t}^{+},y_{t}^{-},x_{t}^{\mathsf{T}})$, so that (3.2) may be written as
@@
$\boldsymbol{\Pi}(y)$ $\coloneqq\begin{bmatrix}\Pi(y)+\Gamma_{1}S(y)&-\Gamma_{1}+\Gamma_{2}&-\Gamma_{2}+\Gamma_{3}&\cdots&-\Gamma_{k-1}\\ S(y)&-I_{p+1}\\ &I_{p+1}&-I_{p+1}\\ &&\ddots&\ddots\\ &&&I_{p+1}&-I_{p+1}\end{bmatrix}$ $\begin{aligned} \boldsymbol{c}&\coloneqq\begin{bmatrix}c\\ 0_{(p+1)(k-1)}\end{bmatrix}\\ \\ \boldsymbol{u}_{t}&\coloneqq\begin{bmatrix}u_{t}\\ 0_{(p+1)(k-1)}\end{bmatrix}\end{aligned}$
so that $\boldsymbol{\Pi}(y)=\boldsymbol{\alpha}\boldsymbol{\beta}(y)^{\mathsf{T}}$, where $\boldsymbol{\alpha}$, $\boldsymbol{\beta}(y)$ and $S(y)$ are as defined in (4.9)–(4.10), we can rewrite the system in companion form as
@@
Our next result is the counterpart of Lemma B.1 for the present case, in that it provides a nonlinear VAR representation for the short-memory components of the model: the equilibrium errors $\xi_{t}\coloneqq\beta(y_{t})^{\mathsf{T}}z_{t}$ and the differences $\Delta z_{t}^{\ast}$, as collected in
@@
Suppose that DGP, CVAR and CO(ii).1 hold. Let
@@
Using the definition of $\boldsymbol{\xi}_{t}$, we may write (B.15) as
@@
Let $\Delta\boldsymbol{\beta}(y_{t})\coloneqq\boldsymbol{\beta}(y_{t})-\boldsymbol{\beta}(y_{t-1})$, and note that $\Delta\boldsymbol{\beta}(y_{t})]^{\mathsf{T}}\boldsymbol{z}_{t}=[\Delta\boldsymbol{\beta}(y_{t})]_{1:1}^{\mathsf{T}}y_{t}$, for $\Delta\boldsymbol{\beta}(y_{t})]_{1:1}$ the first row of $\Delta\boldsymbol{\beta}(y_{t})$, since all its other rows are zero. Thus with the aid of (B.18) we obtain
@@
By construction, for $\delta_{t}$ as defined in (B.16), we have that $y_{t}=\delta_{t}\Delta y_{t}$ whenever $y_{t-1}$ and $y_{t}$ have opposite signs. Since this is also the only case in which $\Delta\boldsymbol{\beta}(y_{t})\neq 0$, it follows that
@@
Substituting this into (B.19) and recalling the definition of $\overline{\boldsymbol{\beta}}_{t}$ thus yields (B.17).
@@
Since $\overline{\boldsymbol{\beta}}_{t}$ is always a convex combination of $\boldsymbol{\beta}(-1)$ and $\boldsymbol{\beta}(+1)$, the autoregressive matrices in the representation (B.17) are contained in the set
@@
By Jungers (2009, Prop. 1.8), $\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A})=\rho_{{\scriptstyle\mathrm{JSR}}}(\{I+\boldsymbol{\beta}(+1)^{\mathsf{T}}\boldsymbol{\alpha},I+\boldsymbol{\beta}(-1)^{\mathsf{T}}\boldsymbol{\alpha}\})$, which is strictly less than unity by CO(ii).2. That $\boldsymbol{\xi}_{t}\sim I^{\ast}(0)$ now follows from ERR, CO(ii).4 and Lemma A.1. In particular, $\xi_{t}=\beta(y_{t})^{\mathsf{T}}z_{t}$, $\Delta y_{t}^{+}$, $\Delta y_{t}^{-}$ and $\Delta x_{t}$ are all $I^{\ast}(0)$, and hence so too is
@@
For the purposes of proving part (ii) of the theorem, we shall initially suppose (as we did in the proof of Theorem 4.1) that DET is replaced by the weaker condition
@@
for $P_{\beta_{\perp}}(+1)$ as defined in (4.7), noting only at the very end of the proof how the result simplifies when DET holds.
Our next step is to extract the common trend components. Recall the definitions of $\beta_{\perp}(y)$ and $\Gamma(1;y)$ in (4.8), and note that we may take
@@
where the r.h.s. is invertible for each $y\in\mathbb{R}$ by CO(ii).3. By Lemma A.1 in Hansen (2005), $\boldsymbol{\beta}(y)^{\mathsf{T}}\boldsymbol{\alpha}$ is therefore also invertible, whence it follows that the complementary projections
@@
are well defined for each $y\in\mathbb{R}$.
Now premultiplying (B.15) by $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}$ and cumulating yields
@@
In view of (B.21) and the definitions of $\boldsymbol{\alpha}_{\perp}$ and $\boldsymbol{\beta}_{\perp}(y)$, the upper left $p\times p$ block of $\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}}(y)$ is $P_{\beta_{\perp}}(y)$, where the latter is as defined in (4.7). Under CO(ii).4, $n^{-1/2}z_{t}\overset{p}{\rightarrow}{\cal Z}_{0}$ for all $t\in\{-k+1,\ldots,0\}$, and so
$n^{-1/2}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{z}_{0}=n^{-1/2}\alpha_{\perp}^{\mathsf{T}}\left[z_{0}-\sum_{i=1}^{k-1}\Gamma_{i}z_{-i}^{\ast}\right]\overset{p}{\rightarrow}\alpha_{\perp}^{\mathsf{T}}\Gamma(1;{\cal Y}_{0}){\cal Z}_{0}.$
Hence, using the result of part (i), the first $p$ rows of (B.22) give